OpenStudy (anonymous):

solve for x. 7x+5=8 2

OpenStudy (texaschic101):

whats that 2 for ?

OpenStudy (anonymous):

7 divided by 2

OpenStudy (texaschic101):

7/2x + 5 = 8 --- is this your problem ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

7/2x + 5 = 8 --- subtract 5 from both sides 7/2x = 8 - 5 7/2x = 3 -- divide both sides by 7/2 x = 3/(7/2) x = 3 * 2/7 x = 6/7 or 7/2x + 5 = 8 -- multiply by the common denominator (2) to get rid of fractions 7x + 10 = 16 -- subtract 10 from both sides 7x = 16 - 10 7x = 6 x = 6/7 either way, you will arrive at the same answer :)

OpenStudy (anonymous):

thankz u so much for helping me